OpenStudy (anonymous):

Find the difference. seven-tenths minus three-tenths one-fifth two-fifths one-half 1

OpenStudy (anonymous):

I think it is 1

OpenStudy (anonymous):

\[\frac{ 7 }{ 10 }- \frac{ 3 }{ 10 }= \frac{ 4 }{ 10 } = \frac{ 2 }{ 5 }\] They have the same denominator, so subtract the numerator and simplify.